# OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models

### v2

@@ 1 Introduction
We believe complex spatial reasoning remains a significant challenge (Gardner, 2011, Baddeley, 1998, Previc, 1998, Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025). Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships (Bar-Anan et al., 2006, Trope & Liberman, 2010, Ramalho et al., 2018). For example, in an emergency, knowing that an AED is “to the right of the door” is insufficient without understanding schematic diagrams, correlating maps with real-world environments, and planning an efficient route. Similarly, tasks like inserting a knife into a rack or flattening a box demand reasoning about object rotation, deformation, and spatial compatibility—far beyond static object placement.
@@ 2.1 Taxonomy of Visual–Spatial Reasoning
Our taxonomy is motivated by two complementary perspectives. (i) Cognitive psychology foundations. Prior research on spatial cognition highlights partly independent faculties—such as visualization, mental rotation, perspective taking, and spatial updating—that can be systematically assessed in humans (Chabris et al., 2006, Meneghetti et al., 2022). These constructs provide a principled scaffold for analyzing how agents perceive, reason about, and act within space. (ii) Moving beyond basic spatial relations. Existing benchmarks are nearly saturated on simple tasks like left–right discrimination, front–back identification, and object counting (Chen et al., 2024, Cai et al., 2024, Du et al., 2024, Szymanska et al., 2024). Yet real-world embodied tasks demand richer reasoning (Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025, Stogiannidis et al., 2025, Lee et al., 2025) about scene dynamics, multi-step logic, physical interaction, and viewpoint transformation.
@@ 2.2 Rationale for Classification
Figure 2: Tasks Demonstration of OmniSpatial. Several representative subtasks are selected for demonstration in each of the four task categories. Note: The questions above are slightly simplified for clarity and conciseness.
@@ 3.1 Overview
We present OmniSpatial, a comprehensive benchmark designed to evaluate vision-language models on spatial reasoning. Rather than pursuing sheer data volume, OmniSpatial emphasizes diversity, structure, and rigor. It now consists of 8.4K carefully curated question–answer pairs, substantially larger than earlier prototypes, and covers a broad spectrum of scenarios that demand reasoning beyond pattern recognition.
@@ 3.2.1 Data Collection
Existing Dataset Images  We integrate two key data sources: MME (Liang et al., 2024) and HOI4D (Liu et al., 2022). MME provides RGB-D data, allowing depth-based spatial inference. We leverage its depth information and manually propose physics-based questions such as “If a red car passes me in 5 seconds, what speed should I maintain?” This ensures realistic distance and motion-based reasoning. HOI4D contains extensive human-object interaction videos. We extract sequential frames to create motion prediction tasks, such as “Where will the hand holding the kettle move next?” By incorporating these datasets, we introduce real-world motion and interaction complexities, further strengthening VLMs’ dynamic reasoning capabilities. Furthermore, to extend our dataset to model training, we partition the dataset into a 1.5K test set and a 6.9K training set. The test set is entirely human-annotated and further integrates samples from several existing datasets. including SpatialViz  (Wang et al., 2025), PhysBench (Chow et al., 2025), ViewSpatial (Li et al., 2025), and DrivingVQA (Corbière et al., 2025), which substantially enhance the diversity of our dataset.
@@ 3.3.2 SpatialCoT: Stimulating Spatial Imagination via Novel Views
Method
Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@
GPT-3.5-turbo (Roumeliotis & Tselikas, 2023) 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
@@
Qwen-VL2.5-3B (Wang et al., 2024b) 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
Qwen-VL2.5-7B (Wang et al., 2024b) 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
Qwen-VL2.5-32B (Wang et al., 2024b) 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
Qwen-VL2.5-72B (Wang et al., 2024b) 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
@@ 4.2 Evaluation Metrics
Method
Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@ 4.3 Main Results
Overall Model Performance  As illustrated in Table˜2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 (OpenAI, 2025b) and Gemini-2.5-pro (Anil et al., 2023), achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and they require a lot of inference time and tokens. (ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B (Zhu et al., 2025) and Qwen-VL2.5-72B (Wang et al., 2024b) achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp. (iii) Specialized Spatial Reasoning Models, due to limitations in dataset coverage and model capacity, struggle to achieve substantial improvements on comprehensive benchmarks.
@@
Impact of PointGraph & Spatial CoT  To test whether structured segmentation improves performance, we apply PointGraph as a preprocessing step for GPT-4.1, Gemini-2.5-flash and Qwen-VL2.5-7B. Results in Table˜3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement. Fig.˜5 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track. Through novel view synthesis facilitated by InstantMesh, both GPT-4.1 and Qwen-VL2.5-7B exhibit significant performance improvements, validating the effectiveness of explicit spatial imagination.
@@ 5.1 Benchmarking Spatial Reasoning
Various studies have introduced innovative benchmarking methodologies (Szymanska et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Qi et al., 2025, Ray et al., 2024) to advance spatial reasoning evaluation. Spatial VQA (Du et al., 2024) was among the first to incorporate spatial information into vision-language models, enabling fundamental spatial relationship reasoning. SpatialBot (Cai et al., 2024) categorized spatial reasoning into various hierarchical levels, extending its applicability to robotic manipulation tasks. RoboSpatial (Song et al., 2024) proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives. VSI-Bench (Yang et al., 2024) combined video data (Chandrasegaran et al., 2024, Liu et al., 2023c, Fu et al., 2024a, Li et al., 2024b, Fang et al., 2024b) with cognitive maps (Momennejad et al., 2023, Apostolopoulos & Groumpos, 2023) to simulate human-like spatial cognition and optimize reasoning in dynamic environments. Recently, SoFar (Qi et al., 2025) introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
@@
While these benchmarks have made significant contributions, a unified framework encompassing a wide range of complex spatial reasoning tasks remains lacking. Inspired by prior spatial reasoning research (Xu et al., 2024b, Wang et al., 2024a, Lin et al., 2014, Nwankwo et al., 2024), we identified several limitations in existing benchmarks, such as reliance on generated images (Szymanska et al., 2024, Fu et al., 2024b, Kamath et al., 2023, Liu et al., 2023a, Rajabi & Kosecka, 2023, Shiri et al., 2024), LLM-generated templates (Linghu et al., 2024, Du et al., 2024, Rädsch et al., 2025), and domain-specific focus (Xie et al., 2025, Chow et al., 2025, Danish et al., 2024). These issues hinder their comprehensiveness and real-world applicability. To address these gaps, our study proposes a comprehensive and integrative spatial reasoning benchmark.
@@ 6 Conclusion
We introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning. OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 8.4 K manually-curated question–answer pairs. Extensive experiments show that state-of-the-art proprietary and open-source VLMs peak at 57% accuracy—over 30 points below human performance—struggling especially with geometric reasoning and non-egocentric perspective taking. To bridge these gaps, we introduce PointGraph for structured scene-graph reasoning and SpatialCoT for viewpoint-aware chain-of-thought, both yielding consistent gains and underscoring the value of structured and multi-view reasoning.
@@ Appendix B Additional Related Works
B.1 Spatial Vision-Language Models
Spatial vision-language models integrate computer vision (He et al., 2022, Oquab et al., 2024, Radford et al., 2021) and natural language processing (Bai et al., 2023, Wang et al., 2024b, Brown et al., 2020, Anil et al., 2023, Touvron et al., 2023a; b, Zhang et al., 2024b) to enhance machine understanding of spatial relationships. Recent research (Qi et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Yuan et al., 2024, Qi et al., 2025) has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments. One of the pioneering works in this domain, SpatialVLM (Chen et al., 2024), has significantly advanced spatial reasoning by constructing an RGB-D based visual question-answering dataset. These models (Hurst et al., 2024; 2024, Reid et al., 2024, Wang et al., 2024b) effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
Further advancements (Liu et al., 2024b; a, Ramakrishnan et al., 2024, Tang et al., 2024, Yang et al., 2023, Rozanova et al., 2021, Wu et al., 2024, Stogiannidis et al., 2025) include SpatialRGPT (Cheng et al., 2024), which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities. Similarly, SpatialBot (Cai et al., 2024) explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments. Recently, SoFar (Qi et al., 2025) proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of VLMs, significantly improving spatial understanding and robotic operation capabilities (Cho et al., 2024, Driess et al., 2023, Mu et al., 2023, Zitkovich et al., 2023, Wake et al., 2024, Huang et al., 2024b, Nasiriany et al., 2024, Duan et al., 2024, Huang et al., 2024a, Qi et al., 2025, He et al., 2025a, Zhang et al., 2025b). However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks. To bridge this gap, our study aims to develop a more comprehensive benchmark that rigorously evaluates spatial reasoning capabilities.
B.2 Spatial Reasoning in Psychology
In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings (Gardner, 2011). To systematically characterize this ability, Buckley et al. (2018)proposed a factor analysis-based framework, distinguishing between spatial visualization, spatial relations (such as mental rotation), and spatial orientation. Malanchini et al. (2020) further identified a strong correlation between spatial orientation and object manipulation skills. Additionally, Newcombe & Shipley (2014)introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic. Furthermore, many studies also have contributed to the development of frameworks (Wai et al., 2009, Lee & Bednarz, 2012, Malanchini et al., 2020, Hegarty, 2010, McGarvey et al., 2018) and evaluation methods (Eliot & Smith, 1983, Uttal et al., 2024; 2013, Hegarty & Waller, 2005) for spatial reasoning. Their framework highlights the distinction between object-centric spatial properties and external reference frames, offering valuable insights for applications such as navigation and path planning. These psychological frameworks provide complementary perspectives for understanding and assessing spatial reasoning, offering theoretical foundations for embodied intelligent systems. Inspired by these classifications, our study proposes a novel taxonomy for visual-spatial reasoning, aiming to advance spatial reasoning research in vision-language models.

### v3

# arXiv:2506.03135v3[cs.CV] 28 Feb 2026

@@ 1 Introduction
We believe complex spatial reasoning remains a significant challenge (Gardner, 2011, Baddeley, 1998, Previc, 1998, Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025a). Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships (Bar-Anan et al., 2006, Trope and Liberman, 2010, Ramalho et al., 2018). For example, in an emergency, knowing that an AED is “to the right of the door” is insufficient without understanding schematic diagrams, correlating maps with real-world environments, and planning an efficient route. Similarly, tasks like inserting a knife into a rack or flattening a box demand reasoning about object rotation, deformation, and spatial compatibility—far beyond static object placement.
@@ 2.1 Taxonomy of Visual–Spatial Reasoning
Our taxonomy is motivated by two complementary perspectives. (i) Cognitive psychology foundations. Prior research on spatial cognition highlights partly independent faculties—such as visualization, mental rotation, perspective taking, and spatial updating—that can be systematically assessed in humans (Chabris et al., 2006, Meneghetti et al., 2022). These constructs provide a principled scaffold for analyzing how agents perceive, reason about, and act within space. (ii) Moving beyond basic spatial relations. Existing benchmarks are nearly saturated on simple tasks like left–right discrimination, front–back identification, and object counting (Chen et al., 2024, Cai et al., 2024, Du et al., 2024, Szymanska et al., 2024). Yet real-world embodied tasks demand richer reasoning (Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025a, Stogiannidis et al., 2025, Lee et al., 2025) about scene dynamics, multi-step logic, physical interaction, and viewpoint transformation.
@@ 2.2 Rationale for Classification
Figure 2: Tasks Demonstration of OmniSpatial. Several representative subtasks are selected for demonstration in each category. Note that the questions above are slightly simplified for clarity.
@@ 3.1 Overview
We present OmniSpatial, a comprehensive benchmark designed to evaluate VLMs on spatial reasoning. Rather than pursuing sheer data volume, OmniSpatial emphasizes diversity, structure, and rigor. It now consists of 8.4K carefully curated QA pairs, substantially larger than earlier prototypes, and covers a broad spectrum of scenarios that demand reasoning beyond pattern recognition.
@@ 3.2.1 Data Collection
Existing Dataset Images  We integrate two key data sources: MME (Liang et al., 2024) and HOI4D (Liu et al., 2022). MME provides RGB-D data, allowing depth-based spatial inference. We leverage its depth information and manually propose physics-based questions such as “If a red car passes me in 5 seconds, what speed should I maintain?” This ensures realistic distance and motion-based reasoning. HOI4D contains extensive human-object interaction videos. We extract sequential frames to create motion prediction tasks, such as “Where will the hand holding the kettle move next?” By incorporating these datasets, we introduce real-world motion and interaction complexities, further strengthening VLMs’ dynamic reasoning capabilities. Furthermore, to extend our dataset to model training, we partition the dataset into a 1.5K test set and a 6.9K training set. The test set is entirely human-annotated and the train set further integrates samples from several existing datasets, including SpatialViz  (Wang et al., 2025), PhysBench (Chow et al., 2025), ViewSpatial (Li et al., 2025), and DrivingVQA (Corbière et al., 2025), which substantially enhance the diversity.
@@ 3.3.2 SpatialCoT: Stimulating Spatial Imagination via Novel Views
Method Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@
GPT-3.5-turbo (Roumeliotis and Tselikas, 2023) 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
@@
Qwen-VL2.5-3B (Wang et al., 2024c) 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
Qwen-VL2.5-7B (Wang et al., 2024c) 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
Qwen-VL2.5-32B (Wang et al., 2024c) 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
Qwen-VL2.5-72B (Wang et al., 2024c) 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
@@ 4.2 Evaluation Metrics
Method Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@ 4.3 Main Results
Overall Model Performance  As illustrated in Table˜2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 (OpenAI, 2025b) and Gemini-2.5-pro (Anil et al., 2023), achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and they require a lot of inference time and tokens. (ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B (Zhu et al., 2025) and Qwen-VL2.5-72B (Wang et al., 2024c) achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp. (iii) Specialized Spatial Reasoning Models, due to limitations in dataset coverage and model capacity, struggle to achieve substantial improvements on comprehensive benchmarks.
@@
Impact of PointGraph & Spatial CoT  To test whether structured segmentation improves performance, we apply PointGraph as a preprocessing step for GPT-4.1, Gemini-2.5-flash and Qwen-VL2.5-7B. Results in Table˜3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement. Fig.˜5 and Table˜4 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track. Through novel view synthesis facilitated by InstantMesh, both GPT-4.1 and Qwen-VL2.5-7B exhibit significant performance improvements, validating the effectiveness of explicit spatial imagination.
@@ 5.1 Benchmarking Spatial Reasoning
Various studies have introduced innovative benchmarking methodologies (Szymanska et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Qi et al., 2025, Ray et al., 2024) to advance spatial reasoning evaluation. Spatial VQA (Du et al., 2024) was among the first to incorporate spatial information into vision-language models, enabling fundamental spatial relationship reasoning. SpatialBot (Cai et al., 2024) categorized spatial reasoning into various hierarchical levels, extending its applicability to robotic manipulation tasks. RoboSpatial (Song et al., 2024) proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives. VSI-Bench (Yang et al., 2024) combined video data (Chandrasegaran et al., 2024, Liu et al., 2023c, Fu et al., 2024a, Li et al., 2024b, Fang et al., 2024b) with cognitive maps (Momennejad et al., 2023, Apostolopoulos and Groumpos, 2023) to simulate human-like spatial cognition and optimize reasoning in dynamic environments. Recently, SoFar (Qi et al., 2025) introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
@@
While these benchmarks have made significant contributions, a unified framework encompassing a wide range of complex spatial reasoning tasks remains lacking. Inspired by prior spatial reasoning research (Xu et al., 2024b, Wang et al., 2024a, Lin et al., 2014, Nwankwo et al., 2024, Wang et al., 2024b), we identified several limitations in existing benchmarks, such as reliance on generated images (Szymanska et al., 2024, Fu et al., 2024b, Kamath et al., 2023, Liu et al., 2023a, Rajabi and Kosecka, 2023, Shiri et al., 2024), LLM-generated templates (Linghu et al., 2024, Du et al., 2024, Rädsch et al., 2025), and domain-specific focus (Xie et al., 2025, Chow et al., 2025, Danish et al., 2024). These issues hinder their comprehensiveness and real-world applicability. To address these gaps, our study proposes a comprehensive and integrative spatial reasoning benchmark.
5.2 Spatial Vision-Language Models
Spatial vision-language models integrate computer vision (He et al., 2022, Oquab et al., 2024, Radford et al., 2021) and natural language processing (Bai et al., 2023, Wang et al., 2024c, Brown et al., 2020, Anil et al., 2023, Touvron et al., 2023a; b, Zhang et al., 2024b) to enhance machine understanding of spatial relationships and mental intelligence. Recent research (Qi et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Yuan et al., 2024, Qi et al., 2025, Chen et al., 2025b) has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments. One of the pioneering works in this domain, SpatialVLM (Chen et al., 2024), has significantly advanced spatial reasoning by constructing an RGBD-based visual question-answering dataset. GS-Reasoner Chen et al. (2025b) introduces 3D grounding as a chain-of-thoughts for spatial reasoning tasks. These models (Hurst et al., 2024, Reid et al., 2024, Wang et al., 2024c) effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
Further advancements (Liu et al., 2024b; a, Ramakrishnan et al., 2024, Tang et al., 2024, Yang et al., 2023, Rozanova et al., 2021, Wu et al., 2024, Stogiannidis et al., 2025) include SpatialRGPT (Cheng et al., 2024), which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities. Similarly, SpatialBot (Cai et al., 2024) explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments. Recently, SoFar (Qi et al., 2025) proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of vision-language models, significantly improving spatial understanding and robotic operation capabilities (Cho et al., 2024, Driess et al., 2023, Mu et al., 2023, Huang et al., 2024b, Nasiriany et al., 2024, Huang et al., 2024a, Qi et al., 2025, He et al., 2025a, Zhang et al., 2025b). However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks. To bridge this gap, our study aims to develop a more comprehensive benchmark that rigorously evaluates spatial reasoning capabilities.
@@ 6 Conclusion
We introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning. OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 8.4 K manually-curated question–answer pairs. Extensive experiments show that state-of-the-art proprietary and open-source VLMs peak at 57% accuracy—over 30 points below human performance—struggling especially with geometric reasoning and non-egocentric perspective taking. To bridge these gaps, we introduce PointGraph for structured scene-graph reasoning and SpatialCoT for viewpoint-aware CoT, both yielding consistent gains and underscoring the value of structured and multi-view reasoning.
Ethic Statement
This work complies with the ICLR Code of Ethics. All datasets are publicly available and used under their respective licenses. The research raises no direct ethical or legal concerns, and the authors are committed to responsible and fair use of the proposed methods.
Reproducibility Statement
We have made every effort to ensure the reproducibility of our work. The proposed model, implementation details, and evaluation protocols are described in detail in the main paper and appendix. All datasets used are publicly available and properly referenced. To further support reproducibility, we release the source code in the supplementary materials.
@@ Appendix B Additional Related Works
B.1 Spatial Reasoning in Psychology
In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings (Gardner, 2011). To systematically characterize this ability, Buckley et al. (2018)proposed a factor analysis-based framework, distinguishing between spatial visualization, spatial relations (such as mental rotation), and spatial orientation. Malanchini et al. (2020) further identified a strong correlation between spatial orientation and object manipulation skills. Additionally, Newcombe and Shipley (2014)introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic. Furthermore, many studies also have contributed to the development of frameworks (Wai et al., 2009, Lee and Bednarz, 2012, Malanchini et al., 2020, Hegarty, 2010, McGarvey et al., 2018) and evaluation methods (Eliot and Smith, 1983, Uttal et al., 2024; 2013, Hegarty and Waller, 2005) for spatial reasoning. Their framework highlights the distinction between object-centric spatial properties and external reference frames, offering valuable insights for applications such as navigation and path planning. These psychological frameworks provide complementary perspectives for understanding and assessing spatial reasoning, offering theoretical foundations for embodied intelligent systems. Inspired by these classifications, our study proposes a novel taxonomy for visual-spatial reasoning, aiming to advance spatial reasoning research in vision-language models.